Saved Help Save & Exit Subn Check my work 9 Problem 5-21 Annuities (LO3) Calculate the present value of the following o-1. Annual payment of $975 for 12 years at 3% interest (Do not round intermediate calculations. Round your answer to 2 decimal places.) 10 polnts eBook Present value 683 85 6-2. Annual payment of S775 for 17 years at 3% interest (Do not round intermediate calculations. Round your answer to 2 decimal places.) Present value a-3. Which option would you preter? k Prev 9ot 15 Next > 23 201 O Type here to search
media%2Fa3d%2Fa3d70b5d-acfb-4e16-a638-47
Answer #1
Annual rate of interest is 3% and 12%
Years PVVF of $ 1 @ 3% Total   PVVF of $ 1 @ 12% Total
0                                   1.00000                          1.00000
1                                   0.97087                          0.89286
2                                   0.94260                          0.79719
3                                   0.91514                          0.71178
4                                   0.88849                          0.63552
5                                   0.86261                          0.56743
6                                   0.83748                          0.50663
7                                   0.81309                          0.45235
8                                   0.78941                          0.40388
9                                   0.76642                          0.36061
10                                   0.74409                          0.32197
11                                   0.72242                          0.28748
12                                   0.70138                                   9.95400                          0.25668                          6.19437
13                                   0.68095                          0.22917
14                                   0.66112                          0.20462
15                                   0.64186                          0.18270
16                                   0.62317                          0.16312
17                                   0.60502                                13.16612                          0.14564                          7.11963
Answer = A-1
Deposite Amount Per Year (A) total PVF of $ 1 in 12 th Year @3% (B) Present Value  
Annual Payment $                                 975.00 $                                   9.954 $                    9,705.15
Answer = A-2
Deposite Amount Per Year (A) total PVF of $ 1 in 17 th Year @3% (B) Present Value  
Annual Payment $                                 775.00 $                                 13.166 $                  10,203.74
Answer = B-1
Deposite Amount Per Year (A) total PVF of $ 1 in 12 th Year @12% (B) Present Value  
Annual Payment $                                 975.00 $                                   6.194 $                    6,039.51
Answer = B-2
Deposite Amount Per Year (A) total PVF of $ 1 in 17 th Year @12% (B) Present Value  
Annual Payment $                                 775.00 $                                   7.120 $                    5,517.71
Answer = B-3
Option 1 = $ 975 a year for 12 year will prefer because it have the greater present value